OpenStudy (anonymous):

2d-9=-29 HELP NEED HOME WORK DONE BY 10 . and have like 20 problems.

OpenStudy (anonymous):

i will help

OpenStudy (anonymous):

please omg.

OpenStudy (anonymous):

d = -10

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

CAN YOU HELP ME MORE?

OpenStudy (anonymous):

http://www.wolframalpha.com will help you with all your math work

OpenStudy (anonymous):

yes i can help

OpenStudy (anonymous):

\[\frac{ m }{ -4 }+5=14\]

OpenStudy (anonymous):

m = -36

OpenStudy (anonymous):

More? andthankkss

OpenStudy (anonymous):

wait that isnt right hold on a sec

OpenStudy (anonymous):

it is cause its an answer

OpenStudy (anonymous):

oh okay im right

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i will help with everything

OpenStudy (anonymous):

-8t-23=-15

OpenStudy (anonymous):

t = -1

OpenStudy (anonymous):

\[\frac{ w }{ 5} + 17 = -3\]

OpenStudy (anonymous):

w = -100

OpenStudy (anonymous):

do you need help with anything else

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

how many more questions?

OpenStudy (anonymous):

but are you sure it is -100

OpenStudy (anonymous):

no im not 100% sure on that one. but im sure on the others

OpenStudy (anonymous):

7x-2=61

OpenStudy (anonymous):

??

OpenStudy (anonymous):

x = 9

OpenStudy (anonymous):

im 100% sure

OpenStudy (anonymous):

can i have a medal and a fan please

OpenStudy (anonymous):

\[\frac{ -a }{ 10}-8=-24\]

OpenStudy (anonymous):

a = 160 (im 80% sure)

OpenStudy (anonymous):

im right

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I;m not doing homework anymore i don't feel like it

OpenStudy (anonymous):

okay message me whenever you need help.

OpenStudy (anonymous):

ok thanxx